# Composing Global Optimizers to Reasoning Tasks via Algebraic Objects in Neural Nets

### v3

@@ Abstract
We prove rich algebraic structures of the solution space for 2-layer neural networks with quadratic activation and $L_{2}$ loss, trained on reasoning tasks in Abelian group (e.g., modular addition). Such a rich structure enables analytical construction of global optimal solutions from partial solutions that only satisfy part of the loss, despite its high nonlinearity. We coin the framework as CoGO (Composing Global Optimizers). Specifically, we show that the weight space over different numbers of hidden nodes of the 2-layer network is equipped with a semi-ring algebraic structure, and the loss function to be optimized consists of monomial potentials, which are ring homomorphisms, allowing partial solutions to be composed into global ones by ring addition and multiplication. Our experiments show that around $95\%$ of the solutions obtained by gradient descent match exactly our theoretical constructions. Although the global optimizers constructed only required a small number of hidden nodes, our analysis on gradient dynamics shows that overparameterization asymptotically decouples training dynamics and is beneficial. We further show that training dynamics favors simpler solutions under weight decay, and thus high-order global optimizers such as perfect memorization are unfavorable. The code is at [^1] https://github.com/facebookresearch/luckmatters/tree/yuandong3/ssl/real-dataset.
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive results in various disciplines (OpenAI, 2024; Anthropic, ; Team, 2024b; a; Dubey et al., 2024; Jiang et al., 2023), while they also make surprising mistakes in basic reasoning tasks (Nezhurina et al., 2024; Berglund et al., 2023). Therefore, it remains an open problem whether it can truly do reasoning tasks. On one hand, existing works demonstrate that the models can learn efficient algorithms (e.g., dynamic programming (Ye et al., 2024) for language structure modeling, etc) and good representations (Jin & Rinard, 2024; Wijmans et al., 2023). Some reports emergent behaviors (Wei et al., 2022) when scaling up with data and model size. On the other hand, many works also show that LLMs cannot self-correct (Huang et al., 2023), and cannot generalize very well beyond the training set for simple tasks (Dziri et al., 2023; Yehudai et al., 2024; Ouellette et al., 2023), let alone complicated planning tasks (Kambhampati et al., 2024; Xie et al., 2024).
@@
In this work, we systematically analyze 2-layer neural networks with quadratic activation and $L_{2}$ loss on predicting the outcome of group multiplication in Abelian group $G$, which is an extension of modular addition. We find that global optimizers can be constructed algebraically from small partial solutions that are optimal only for parts of the loss. We achieve this by showing that (1) for the 2-layer network, there exists a semi-ring structure over the weights space across different order (i.e., number of hidden nodes or network width), with specifically defined addition and multiplication (Sec. 4.1), and (2) the $L_{2}$ loss is a function of monomial potentials (MPs), which are ring homomorphisms (Theorem 1) that allow compositions of partial solutions into global ones using ring addition and multiplication.
As a result, our theoretical framework, named CoGO (i.e., Composing Global Optimizers), successfully constructs two distinct types of Fourier-based global optimizers of per-frequency order 4 (or “$2\times 2$”) and order 6 (or “$2\times 3$”), and a global optimizer of order $d^{2}$ that correspond to perfect memorization. Empirically, we demonstrate that around $95\%$ of the solutions obtained from gradient descent (with weight decay) have the predicted structure and match exactly with our theoretical construction of order-4 and order-6 solutions. In addition, we also analyze the training dynamics, and show that the dynamics favors low-order global optimizers, since global optimizers algebraically connected by ring multiplication can be proven to also be topologically connected. Therefore, high-order solution like perfect memorization is unfavorable in the dynamics. When the network width goes to infinity, the dynamics of monomial potentials becomes decoupled, demystifying why overparameterization improves the performance.
@@ 2 Related Works
Figure 1: Overview of proposed theoretical framework CoGO. (1) The family of 2-layer neural networks, $\mathcal{Z}$, form a semi-ring algebraic structure (Theorem 2) with ring addition and multiplication (Def. 5). $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$ where $\mathcal{Z}_{q}$ is a collection of all weights with order-$q$ (i.e., $q$ hidden nodes). (2) For outcome prediction of Abelian group multiplication, the MSE loss $\ell({\bm{z}})$ is a function of monomial potentials (MPs) $r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$ (Theorem 1), which are ring homomorphisms (Theorem 3). (3) Thanks to the property of ring homomorphism, global optimizers to MSE loss $\ell({\bm{z}})$ with quadratic activation can be constructed algebraically from partial solutions that only satisfy a subset of constraints (Sec. 5.1) using ring addition and multiplication, instead of running gradient descent. Examples include Fourier solution ${\bm{z}}_{F6}$ (Corollary 2) and ${\bm{z}}_{F4/6}$ (Corollary 4) and perfect memorization solution ${\bm{z}}_{M}$ (Corollary 3). In Sec. 6, we analyze the role played of MPs in gradient dynamics, showing that the dynamics favors low-order global optimizers (Theorem 6) under weight decay regularization, and the dynamics of MPs become decoupled with infinite width (Theorem 7). In Sec. 7 we show that the gradient descent solutions match exactly with our theoretical construction.
3 Decoupling $L_{2}$ Loss for reasoning tasks of Abelian group
@@
Basic Ring theory. A set $\mathcal{Z}$ forms a ring, if there exists two operations, addition $+$ and multiplication *, so that (1) $\langle\mathcal{Z},+\rangle$ forms an Abelian group, (2) $\langle\mathcal{Z},*\rangle$ is a monoid (i.e., a group without inverse), and (3) multiplication distributes with addition (i.e., $a*(b+c)=a*b+a*c$ and $(b+c)*a=b*a+c*a$). $\mathcal{Z}$ is called a semi-ring if $\langle\mathcal{Z},+\rangle$ is a monoid.
@@
$\ell=\sum_{i}\Big{\|}P^{\perp}_{1}\left(\frac{1}{2d}{\bm{o}}[i]-{\bm{e}}_{l[i]}\right)\Big{\|}^{2},\quad\quad{\bm{o}}[i]=\sum_{j}\mathbf{w}_{cj}\sigma(\mathbf{w}_{aj}^{\top}{\bm{e}}_{g_{1}[i]}+\mathbf{w}_{bj}^{\top}{\bm{e}}_{g_{2}[i]})$ (1)
@@
Figure 2: Problem setup. For group multiplication $l[i]=g_{1}[i]g_{2}[i]\mod d$, we use a 2-layer network with quadratic activation to predict $l[i]$ from $g_{1}[i]$ and $g_{2}[i]$. See Eqn. 1 for the loss function.
Architectures. In Eqn. 1, we use quadratic activation $\sigma(x)=x^{2}$ (Du & Lee, 2018; Allen-Zhu & Li, 2023), $P_{1}^{\perp}=I-\frac{1}{d}{\bm{1}}{\bm{1}}^{\top}$ is the zero-mean projection, $\mathbf{w}_{aj},\mathbf{w}_{bj},\mathbf{w}_{cj}\in\mathbb{R}^{d}$ are learnable parameters ($1\leq j\leq q$). See Fig. 2. Note that variants of quadratic activation have been used empirically, e.g. squared ReLU and gated activations (So et al., 2021; Shazeer, 2020; Zhang et al., 2024).
We can extend our framework to group action prediction, in which $g_{2}$ may not be a group element but any object (e.g., a discrete state in reinforcement learning). See Appendix E for more details.
Let $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}\in\mathbb{C}^{d}$ be the scaled Fourier bases (or more formally, character function of the finite Abelian group $G$, see Appendix A). Then the weight vector $\mathcal{W}:=\{\mathbf{w}_{j}\}$ can be written as:
where ${\bm{z}}:=\{z_{pkj}\}$ are the complex coefficients, $p\in\{a,b,c\}$, $0\leq k<d$ and $j$ runs through $q$ hidden nodes. For convenience, we define $\bm{\phi}_{-k}:=\overline{\bm{\phi}}_{k}$ as the (complex) conjugate representation of $\bm{\phi}_{k}$. We exclude $\bm{\phi}_{0}\equiv 1$ because the constant bias term has been filtered out by the top-down gradient from the loss function. Leveraging the property of quadratic activation functions, we can write down the loss function analytically (see Appendix A):
@@
[…]
@@
Eqn. 3 contains different $r$ terms, which play an important role in determining global optimizers.
@@
Let $R:=\{r\}$ be a collection of $r$ terms. The weight ${\bm{z}}$ is said to have 00-set $R_{0}$ and $1$-set $R_{1}$ (or 0/1-sets $(R_{0},R_{1})$), if $r({\bm{z}})=0$ for all $r\in R_{0}$ and $r({\bm{z}})=1$ for all $r\in R_{1}$.
With 0/1-sets, we can characterize rough structures of the global optimizers to the loss:
Lemma 1 (A Sufficient Conditions of Global optimizers of Eqn. 3).
@@
then it is a global optimizer with zero loss $\ell({\bm{z}})=0$. Here $R_{\mathrm{g}}:=\{r_{kkk},k\neq 0\}$, $R_{\mathrm{c}}:=\{r_{k_{1}k_{2}k},k_{1},k_{2},k\mathrm{\ not\ all\ equal}\}$, $R_{\mathrm{n}}:=\{r_{p,k^{\prime},-k^{\prime},k}\}$ and $R_{*}:=\{r_{p,k^{\prime},m-k^{\prime},k},m\neq 0\}$.
Lemma 1 provides sufficient conditions since there may exist solutions that achieve global optimum (e.g., $\sum_{k^{\prime}}r_{p,k^{\prime},m-k^{\prime},k}({\bm{z}})=0$ but $r_{p,k^{\prime},m-k^{\prime},k}({\bm{z}})\neq 0$). However, as we will see, it already leads to rich algebraic structure, and serves as a good starting point. Directly finding the global optimizers using Eqn. 4 can be a bit complicated and highly non-intuitive, due to highly nonlinear structure of Eqn. 3. However, there are nice structures we can leverage, as we will demonstrate below.
4 Beyond Fixed Parameter Space: The Semi-ring structure
4.1 The semi-ring structure of the solution space
We define the weight space $\mathcal{Z}_{q}=\{{\bm{z}}\}$ to include all the weight matrices with $q$ hidden nodes ($\mathcal{Z}_{0}$ means an empty network), and $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$ be the solution space of all different number of hidden nodes. Interestingly, $\mathcal{Z}$ naturally is equipped with a semi-ring structure, and each term of the loss function can effective interact with such a semi-ring structure, yielding provable global optimizers, including both the Fourier solutions empirically reported in previous works (Zhou et al., 2024; Gromov, 2023), and the perfect memorization solution (Morwani et al., 2023).
To make our argument formal, we start with a few definitions.
@@ 4.1 The semi-ring structure of the solution space
$\alpha{\bm{z}}\in\mathcal{Z}$ is element-wise multiplication $[\alpha z_{pkj}]$ of ${\bm{z}}\in\mathcal{Z}$.
@@
For any two solutions ${\bm{z}}_{1}:=\{z^{(1)}_{pkj}\}$ and ${\bm{z}}_{2}:=\{z^{(2)}_{pkj}\}$, we can define their operations:
@@
As we will see, the semi-ring structure of $\mathcal{Z}$ paves the way to construct explicitly global optimizers.
4.2 The Monomial Potentials and its connection to semi-ring $\mathcal{Z}$
Now let us study the structure of the loss function Eqn. 3 and how they are related to the semi-ring structure of $\mathcal{Z}$. For this, we first define the concept of monomial potentials:
Definition 6 (Monomial potential (MP)).
$r({\bm{z}}):=\sum_{j}\prod_{{p,k}\in\mathrm{idx}(r)}z_{pkj}$ is called monomial potential (MP), where $\mathrm{idx}(r)$ specifies the indices involved in the monomial terms.
Following this definition, terms in the loss function (Theorem 1) are examples of MPs.
Observation 1 (Specific MPs).
$r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$ defined in Theorem 1 are MPs.
So what is the relationship between MPs, which are functions that map a weight ${\bm{z}}$ to a complex scalar, and the semi-ring structure of $\mathcal{Z}$? The following theorem tells that MPs are ring homomorphisms, that is, these mappings respect addition and multiplication:
@@ 4.2 The Monomial Potentials and its connection to semi-ring $\mathcal{Z}$
For any monomial potential $r:\mathcal{Z}\mapsto\mathbb{C}$, $r({\bm{1}})=1$, $r({\bm{z}}_{1}+{\bm{z}}_{2})=r({\bm{z}}_{1})+r({\bm{z}}_{2})$ and $r({\bm{z}}_{1}*{\bm{z}}_{2})=r({\bm{z}}_{1})r({\bm{z}}_{2})$ and thus $r$ is a ring homomorphism.
@@
Since the loss function $\ell({\bm{z}})$ depends on the weight ${\bm{z}}$ entirely through $r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$, which are MPs, due to the property of ring homomorphism, it is possible to construct a global optimizer from partial solutions that satisfy only some of the constraints [^2] Mathematically, the kernel $\mathrm{Ker}(r):=\{{\bm{z}}:r({\bm{z}})=0\}$ of a ring homomorphism $r$ is an ideal of the ring, and the intersection of ideals are still ideals. For brevity, we omit the formal definitions.:
@@
If ${\bm{z}}_{1}$ has 0/1-sets $(R_{1}^{-},R^{+}_{1})$ and ${\bm{z}}_{2}$ has 0/1-sets $(R_{2}^{-},R^{+}_{2})$, then (1) ${\bm{z}}_{1}*{\bm{z}}_{2}$ has 0/1-sets $(R_{1}^{-}\cup R_{2}^{-},R^{+}_{1}\cap R^{+}_{2})$. (2) ${\bm{z}}_{1}+{\bm{z}}_{2}$ have 0/1-sets $(R_{1}^{-}\cap R_{2}^{-},R^{+}_{1}\cup R^{+}_{2})$.
Once we reach 0/1-sets $(R_{\mathrm{c}}\cup R_{\mathrm{n}}\cup R_{*},R_{\mathrm{g}})$, we find a global optimizer. In addition, we also immediately know that there exists infinitely many global optimizers, via ring multiplication (Def. 5):
@@
If ${\bm{z}}$ is a global optimizer and ${\bm{y}}$ is a unit, then ${\bm{z}}*{\bm{y}}$ is also a global optimizer.
5 Composing Global Optimizers
5.1 Constructing Partial Solutions with Polynomials
While intuitively one can get global optimizers by manually crafting some partial solutions and combining, in this section, we provide a more systematic approach to compose global optimizers as follows. Since $\mathcal{Z}$ enjoys a semi-ring structure, we consider a polynomial in $\mathcal{Z}$ in the following form:

### v4

# Composing Global Solutions to Reasoning Tasks via Algebraic Objects in Neural Nets

@@ Abstract
We prove rich algebraic structures of the solution space for 2-layer neural networks with quadratic activation and $L_{2}$ loss, trained on reasoning tasks in Abelian group (e.g., modular addition). Such a rich structure enables analytical construction of global optimal solutions from partial solutions that only satisfy part of the loss, despite its high nonlinearity. We coin the framework as CoGS (Composing Global Solutions). Specifically, we show that the weight space over different numbers of hidden nodes of the 2-layer network is equipped with a semi-ring algebraic structure, and the loss function to be optimized consists of sum potentials, which are ring homomorphisms, allowing partial solutions to be composed into global ones by ring addition and multiplication. Our experiments show that around $95\%$ of the solutions obtained by gradient descent match exactly our theoretical constructions. Although the global solutions constructed only required a small number of hidden nodes, our analysis on gradient dynamics shows that overparameterization asymptotically decouples training dynamics and is beneficial. We further show that training dynamics favors simpler solutions under weight decay, and thus high-order global solutions such as perfect memorization are unfavorable. The code is open sourced [^1] https://github.com/facebookresearch/luckmatters/tree/yuandong3/ssl/real-dataset.
@@ 1 Introduction
Large Language Models (LLMs) have shown impressive results in various disciplines (OpenAI, 2024; Anthropic, ; Team, 2024b, a; Dubey et al., 2024; Jiang et al., 2023), while they also make surprising mistakes in basic reasoning tasks (Nezhurina et al., 2024; Berglund et al., 2023). Therefore, it remains an open problem whether it can truly do reasoning tasks. On one hand, existing works demonstrate that the models can learn efficient algorithms (e.g., dynamic programming (Ye et al., 2024) for language structure modeling, etc) and good representations (Jin & Rinard, 2024; Wijmans et al., 2023). Some reports emergent behaviors (Wei et al., 2022) when scaling up with data and model size. On the other hand, many works also show that LLMs cannot self-correct (Huang et al., 2023), and cannot generalize very well beyond the training set for simple tasks (Dziri et al., 2023; Yehudai et al., 2024; Ouellette et al., 2023), let alone complicated planning tasks (Kambhampati et al., 2024; Xie et al., 2024).
@@
In this work, we systematically analyze 2-layer neural networks with quadratic activation and $L_{2}$ loss on predicting the outcome of group multiplication in Abelian group $G$, which is an extension of modular addition. We find that global solutions can be constructed algebraically from small partial solutions that are optimal only for parts of the loss. We achieve this by showing that (1) for the 2-layer network, there exists a semi-ring structure over the weights space across different order (i.e., number of hidden nodes or network width), with specifically defined addition and multiplication (Sec. 4), and (2) the $L_{2}$ loss is a function of sum potentials (SPs), which are ring homomorphisms (Theorem 1) that allow compositions of partial solutions into global ones using ring operations.
As a result, our theoretical framework, named CoGS (i.e., Composing Global Solutions), successfully constructs two distinct types of Fourier-based global solutions of per-frequency order 4 (or $2\times 2$”) and order 6 (or $2\times 3$”), and a global solution of order $d^{2}$ that correspond to perfect memorization. Empirically, we demonstrate that around $95\%$ of the solutions obtained from gradient descent (with weight decay) have the predicted structure and match exactly with our theoretical construction of order-4 and order-6 solutions. In addition, we also analyze the training dynamics, and show that the dynamics favors low-order global solutions, since global solutions algebraically connected by ring multiplication can be proven to also be topologically connected. Therefore, high-order solution like perfect memorization is unfavorable in the dynamics. When the network width goes to infinity, the dynamics of sum potentials becomes decoupled, demystifying why overparameterization improves the performance.
@@ 2 Related Works
Figure 1: Overview of proposed theoretical framework CoGS. (1) The family of 2-layer neural networks, $\mathcal{Z}$, form a semi-ring algebraic structure (Theorem 2) with ring addition and multiplication (Def. 5). $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$ where $\mathcal{Z}_{q}$ is a collection of all weights with order-$q$ (i.e., $q$ hidden nodes). (2) For outcome prediction of Abelian group multiplication, the MSE loss $\ell({\bm{z}})$ is a function of sum potentials (SPs) $r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$ (Theorem 1), which are ring homomorphisms (Theorem 3). (3) Thanks to the property of ring homomorphism, global solutions to MSE loss $\ell({\bm{z}})$ with quadratic activation can be constructed algebraically from partial solutions that only satisfy a subset of constraints (Sec. 5) using ring addition and multiplication, instead of running gradient descent. Examples include Fourier solution ${\bm{z}}_{F6}$ (Corollary 2) and ${\bm{z}}_{F4/6}$ (Corollary 4) and perfect memorization solution ${\bm{z}}_{M}$ (Corollary 3). In Sec. 6, we analyze the role played of SPs in gradient dynamics, showing that the dynamics favors low-order global solutions (Theorem 5) under weight decay regularization, and the dynamics of SPs become decoupled with infinite width (Theorem 6). In Sec. 7 we show that the gradient descent solutions match exactly with our theoretical construction.
3 $L_{2}$ Loss Decoupling for Abelian group
@@
Basic Ring theory. A set $\mathcal{Z}$ forms a ring, if there exists two operations, addition ++ and multiplication *, so that (1) $\langle\mathcal{Z},+\rangle$ forms an Abelian group, (2) $\langle\mathcal{Z},*\rangle$ is a monoid (i.e., a group without inverse), and (3) multiplication distributes with addition (i.e., a(b+c)=ab+aca*(b+c)=a*b+a*c and (b+c)a=ba+ca(b+c)*a=b*a+c*a). $\mathcal{Z}$ is called a semi-ring if $\langle\mathcal{Z},+\rangle$ is a monoid.
@@
$\ell=\sum_{i}\Big\|P^{\perp}_{1}\left(\frac{1}{2d}{\bm{o}}[i]-{\bm{e}}_{l[i]}\right)\Big\|^{2},\quad{\bm{o}}[i]=\sum_{j}\mathbf{w}_{cj}\sigma(\mathbf{w}_{aj}^{\top}{\bm{e}}_{g_{1}[i]}+\mathbf{w}_{bj}^{\top}{\bm{e}}_{g_{2}[i]})$ (1)
@@
Architectures. In Eqn. 1, we use quadratic activation $\sigma(x)=x^{2}$ (Du & Lee, 2018; Allen-Zhu & Li, 2023), $P_{1}^{\perp}=I-\frac{1}{d}{\bm{1}}{\bm{1}}^{\top}$ is the zero-mean projection, $\mathbf{w}_{aj},\mathbf{w}_{bj},\mathbf{w}_{cj}\in\mathbb{R}^{d}$ are learnable parameters ($1\leq j\leq q$). Note that variants of quadratic activation have been used empirically, e.g. squared ReLU and gated activations (So et al., 2021; Shazeer, 2020; Zhang et al., 2024).
We can extend our framework to group action prediction, in which $g_{2}$ may not be a group element but any object (e.g., a discrete state in reinforcement learning), See Appendix F.
Let $\bm{\phi}_{k}=[\phi_{k}(g)]_{g\in G}\in\mathbb{C}^{d}$ be the scaled Fourier bases (or more formally, character function of the finite Abelian group $G$, see Appendix B). Then the weight vector $\mathcal{W}:=\{\mathbf{w}_{j}\}$ can be written as:
$\mathbf{w}_{aj}=\sum_{k\neq 0}z_{akj}\bm{\phi}_{k},\quad\mathbf{w}_{bj}=\sum_{k\neq 0}z_{bkj}\bm{\phi}_{k},\quad\mathbf{w}_{cj}=\sum_{k\neq 0}z_{ckj}\bar{\bm{\phi}}_{k}$ (2)
where ${\bm{z}}:=\{z_{pkj}\}$ are the complex coefficients, $p\in\{a,b,c\}$, $0\leq k<d$ and $j$ runs through $q$ hidden nodes. For convenience, we define $\bm{\phi}_{-k}:=\overline{\bm{\phi}}_{k}$ as the (complex) conjugate representation of $\bm{\phi}_{k}$. We exclude $\bm{\phi}_{0}\equiv 1$ because the constant bias term has been filtered out by the top-down gradient from the loss function. Leveraging the property of quadratic activation functions, we can write down the loss function analytically (see Appendix B):
@@
$\ell_{k}=-2r_{kkk}\!+\!\sum_{k_{1}k_{2}}|r_{k_{1}k_{2}k}|^{2}+\frac{1}{4}\Big|\sum_{p\in\{a,b\}}\sum_{k^{\prime}}r_{p,k^{\prime},-k^{\prime},k}\Big|^{2}+\frac{1}{4}\sum_{m\neq 0}\sum_{p\in\{a,b\}}\Big|\sum_{k^{\prime}}r_{p,k^{\prime},m-k^{\prime},k}\Big|^{2}$ (3)
@@
Eqn. 3 contains different $r$ terms, which play an important role in determining global solutions.
@@
Let $R:=\{r\}$ be a collection of $r$ terms. The weight $\bm{z}$} is said to have 0-set $R_{0}$ and $1$-set $R_{1}$ (or 0/1-sets $(R_{0},R_{1})$), if $r({\bm{z}})=0$ for all $r\in R_{0}$ and $r({\bm{z}})=1$ for all $r\in R_{1}$.
With 0/1-sets, we can characterize rough structures of the global solutions to the loss:
Lemma 1 (A Sufficient Conditions of Global Solutions of Eqn. 3).
@@
then it is a global solution with $\ell({\bm{z}})\!=\!0$. Here $R_{\mathrm{g}}\!:=\!\{r_{kkk},k\neq 0\}$, $R_{\mathrm{c}}:=\{r_{k_{1}k_{2}k},k_{1},k_{2},k\mathrm{\ not\ all\ equal}\}$, $R_{\mathrm{n}}:=\{r_{p,k^{\prime},-k^{\prime},k}\}$ and $R_{*}:=\{r_{p,k^{\prime},m-k^{\prime},k},m\neq 0\}$.
Lemma 1 provides sufficient conditions since there may exist solutions that achieve global optimum (e.g., $\sum_{k^{\prime}}r_{p,k^{\prime},m-k^{\prime},k}({\bm{z}})=0$ but $r_{p,k^{\prime},m-k^{\prime},k}({\bm{z}})\neq 0$). However, as we will see, it already leads to rich algebraic structure, and serves as a good starting point. Directly finding the global solutions using Eqn. 4 can be a bit complicated and highly non-intuitive, due to highly nonlinear structure of Eqn. 3. However, there are nice structures we can leverage, as we will demonstrate below.
4 Algebraic Property of the Weight Space
We define the weight space $\mathcal{Z}_{q}=\{{\bm{z}}\}$ to include all the weight matrices with $q$ hidden nodes ($\mathcal{Z}_{0}$ means an empty network), and $\mathcal{Z}=\bigcup_{q\geq 0}\mathcal{Z}_{q}$ be the solution space of all different number of hidden nodes. Interestingly, $\mathcal{Z}$ naturally is equipped with a semi-ring structure, and each term of the loss function can effective interact with such a semi-ring structure, yielding provable global solutions, including both the Fourier solutions empirically reported in previous works (Zhou et al., 2024; Gromov, 2023), and the perfect memorization solution (Morwani et al., 2023).
To formalize our argument, we start with a few definitions.
@@ 4 Algebraic Property of the Weight Space
$\alpha{\bm{z}}\in\mathcal{Z}$ is element-wise multiplication $\alpha z_{pkj}]$ of ${\bm{z}}\in\mathcal{Z}$.
@@
We define operations for ${\bm{z}}_{1}:=\{z^{(1)}_{pkj}\}$ and ${\bm{z}}_{2}:=\{z^{(2)}_{pkj}\}$:
@@
As we shall see, Thm. 2 allows the construction of global solutions. Now let us explore the structure of the loss (Eqn. 3), which turns out to be connected with the semi-ring structure of $\mathcal{Z}$. For this, we first define the concept of sum potentials:
Definition 6 (Sum potential (SP)).
Let sum potential be $r({\bm{z}}):=\sum_{j}\prod_{{p,k}\in\mathrm{idx}(r)}z_{pkj}$ which takes the summation of a monomial term over all hidden nodes. Here $\mathrm{idx}(r)$ specifies the terms involved.
Following this definition, terms in the loss function (Theorem 1) are examples of SPs.
Observation 1 (Specific SPs).
$r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$ defined in Theorem 1 are SPs.
So what is the relationship between SPs, which are functions that map a weight $\bm{z}$} to a complex scalar, and the semi-ring structure of $\mathcal{Z}$? The following theorem tells that SPs are ring homomorphisms, that is, these mappings respect addition and multiplication:
@@
For any sum potential $r:\mathcal{Z}\mapsto\mathbb{C}$, $r({\bm{1}})=1$, $r({\bm{z}}_{1}+{\bm{z}}_{2})=r({\bm{z}}_{1})+r({\bm{z}}_{2})$ and $r({\bm{z}}_{1}*{\bm{z}}_{2})=r({\bm{z}}_{1})r({\bm{z}}_{2})$ and thus $r$ is a ring homomorphism.
@@
Since the loss function $\ell({\bm{z}})$ depends on the weight $\bm{z}$} entirely through $r_{k_{1}k_{2}k}({\bm{z}})$ and $r_{pk_{1}k_{2}k}({\bm{z}})$, which are SPs, due to the property of ring homomorphism, it is possible to construct a global solution from partial solutions that satisfy only some of the constraints [^2] Mathematically, the kernel $\mathrm{Ker}(r):=\{{\bm{z}}:r({\bm{z}})=0\}$ of a ring homomorphism $r$ is an ideal of the ring, and the intersection of ideals are still ideals. For brevity, we omit the formal definitions.:
@@
[…]
Once we reach 0/1-sets $(R_{\mathrm{c}}\cup R_{\mathrm{n}}\cup R_{*},R_{\mathrm{g}})$, we find a global solution. In addition, we also immediately know that there exists infinitely many global solutions, via ring multiplication (Def. 5):
@@
If $\bm{z}$} is a global solution and $\bm{y}$} is a unit, then ${\bm{z}}*{\bm{y}}$ is also a global solution.
5 Composing Global Solutions
Constructing Partial Solutions with Polynomials. While intuitively one can get global solutions by manually crafting some partial solutions and combining, in this section, we provide a more systematic approach to compose global solutions as follows. Since $\mathcal{Z}$ enjoys a semi-ring structure, we consider a polynomial in $\mathcal{Z}$: